An electric kettle used to prepare tea, takes 2 minutes to boil 4 cups of water (1 cup contains 200 cc of water) if the room temperature is 25°C. (a) If the cost of power consumption is Rs. 1.00 per unit (1 unit – 1000 watt-hour), calculate the coat of coiling 4 cups of water. (b) What will be the corresponding cost if the room temperature drops to 5°C?

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. Volume of water boiled = 4 * 200 cc = 800 cc T base 1 = 25°C T base 2 = 100°C ⇒ T base 2 - T base 1 = 75°C Mass of water boiled = 800 * 1 = 800 gm = 0.8 kg Q(heat req.) = MS∆θ = 0.8 * 4200 * 75 = 252000 J. 1000 watt – hour = 1000 * 3600 watt-sec = 1000* 3600 J No. of units = 252000/1000 * 3600 = 0.07 = 7 paise (b) Q = mS∆T = 0.8 * 4200 * 95 J No. of units = 0.8 * 4200 *95/1000 *3600 = 0.0886 = 0.09 Money consumed = 0.09 Rs = 9 paise.